# SemFlow: Binding Semantic Segmentation and Image Synthesis via Rectified Flow

### v1

@@ 4.3 Ablation Studies
Figure 4: Semantic image synthesis results on FFHQ dataset. Semantic masks are colored to show different semantic components. SemFlow w/ Perturbation indicates the finite perturbation operation in Eq. 14.
Perturbation. Fig. 4 shows the synthesized results of SemFlow on the FFHQ [27] dataset and demonstrates the importance of perturbation operation. First, perturbation enables our approach to generate multi-modal results from one semantic layout using different noises. The synthesized results show diverse appearances, such as skin, hair color, and the background, and exhibit good consistency with the masks. Moreover, we also show the synthesis results of the model without perturbation in Fig. 4. It is witnessed that the model without perturbation is only able to generate uni-modal results, and the synthesized images exhibit a loss of detail and suffer from severe over-smoothing problems.
@@
Tab. 3 compares the performance of DSM and SemFlow on segmentation tasks under different inference steps, where DSM sampled with DDPM and DDIM represent SDE and ODE modeling, respectively. Overall, SemFlow outperforms DSM by a remarkable margin regardless of the inference steps. Specifically, the DSMs’ performance declines markedly as the number of steps decreases. With 5 steps, the DSM-DDPM achieves a mIoU of 10.1, while the DSM-DDIM achieves 7.5 mIoU. On the contrary, our SemFlow achieves 35.9 even with one step.
Fig. 5 provides synthesis results of FFHQ with different inference steps. Although fewer steps will result in some loss of detail and make the images appear smoother, the results of one-step synthesis are still competitive. This indicates that the trajectory between the two distributions is straightened. Fig.6 and Fig. 7 visualize the latent variable sampled from the trajectory. The transition between $z_{0}$ and $z_{1}$ is smooth.
Table 3: Semantic segmentation results with different inference steps on COCO dataset. mIoU is used as the metric.
@@
DSM DDIM - 0.2 7.5 15.6 20.4 23.1 24.2 22.9
DDPM - 0.2 10.1 19.5 26.1 27.5 28.8 29.3
SemFlow Euler 35.9 41.5 48.0 49.2 48.7 48.2 47.9 46.8
Figure 5: Synthesis results with different inference steps on FFHQ dataset. We use the forward Euler method to get numerical solutions. Our approach obtains competitive results even with only one inference step.
Figure 6: Visualization of latent variables on the trajectory from $z_{1}$ to $z_{0}$ (Semantic image synthesis). Top row: COCO. Bottom row: Cityscapes.
Figure 7: Visualization of latent variables on the trajectory from $z_{0}$ to $z_{1}$ (Semantic segmentation). Top row: COCO. Bottom row: Cityscapes.
@@ 5 Conclusion
Acknowledgement. This work is supported by the National Key Research and Development Program of China (No. 2023YFC3807600).

### v2

@@ 4.3 Ablation Studies
Figure 4: Semantic image synthesis results on CelebAMask-HQ dataset. Semantic masks are colored to show different semantic components. SemFlow w/ Perturbation indicates the finite perturbation operation in Eq. 14.
Perturbation. Fig. 4 shows the synthesized results of SemFlow on the CelebAMask-HQ dataset and demonstrates the importance of perturbation operation. First, perturbation enables our approach to generate multi-modal results from one semantic layout using different noises. The synthesized results show diverse appearances, such as skin, hair color, and the background, and exhibit good consistency with the masks. Moreover, we also show the synthesis results of the model without perturbation in Fig. 4. It is witnessed that the model without perturbation is only able to generate uni-modal results, and the synthesized images exhibit a loss of detail and suffer from severe over-smoothing problems.
@@
Tab. 2 compares the performance of DSM and SemFlow on segmentation tasks under different inference steps, where DSM sampled with DDPM and DDIM represent SDE and ODE modeling, respectively. Overall, SemFlow outperforms DSM by a remarkable margin regardless of the inference steps. Specifically, the DSMs’ performance declines markedly as the number of steps decreases. With 5 steps, the DSM-DDPM achieves a mIoU of 12.3, while the DSM-DDIM achieves 9.5 mIoU. On the contrary, our SemFlow achieves 28.3 even with one step.
Fig. 5 provides semantic image synthesis results with different inference steps. Although fewer steps will result in some loss of detail and make the images appear smoother, the results of one-step synthesis are still competitive. This indicates that the trajectory between the two distributions is straightened. Fig. 6 and Fig. 7 visualizes the latent variable sampled from the trajectory. The transition between $z_{0}$ and $z_{1}$ is smooth.
Table 2: Semantic segmentation results with different inference steps on COCO-Stuff dataset. mIoU is used as the metric.
@@
DSM DDIM - 0.1 4.0 9.5 13.6 14.9 15.7 16.1
DDPM - 0.1 5.3 12.3 17.5 18.9 19.6 20.2
SemFlow Euler 28.3 31.0 36.9 38.4 38.6 38.4 38.3 38.3
Figure 5: Image synthesis results with different inference steps. We use the forward Euler method to get numerical solutions. Our approach obtains competitive results even with only one inference step.
Figure 6: Visualization of latent variables on the trajectory from $z_{1}$ to $z_{0}$ (Semantic image synthesis). Top row: COCO-Stuff. Bottom row: Cityscapes.
Figure 7: Visualization of latent variables on the trajectory from $z_{0}$ to $z_{1}$ (Semantic segmentation). Top row: COCO-Stuff. Bottom row: Cityscapes.
@@ 5 Conclusion
Acknowledgement. This work was supported by the National Key Research and Development Program of China (No. 2023YFC3807600).
Appendix A More Details on Method
Amplitude of Perturbation. In the main paper, the perturbation follows the uniform distribution $\epsilon\sim\mathrm{U}(-\beta,\beta)$. We set $20<\beta<s/$ 2 to prevent the semantic labels from changing. Here is the proof.
Let $M=(m_{0},m_{1},m_{2})$ be the anchor for category $c_{1}$, $M^{\prime}=M+\epsilon$ be the anchor with perturbation, and $P=(p_{0},p_{1},p_{2})$ be another anchor for category $c_{2},c_{1}\neq c_{2}$. First, it is obvious that $E(M^{\prime})=E(M)=M$, which means that adding perturbation will not change the estimation of original anchors. Then we only need to prove $||M^{\prime}-M||_{2}<||M^{\prime}-P||_{2}$, which is equivalent to $\sum_{i\in[3]}\epsilon_{i}^{2}<\sum_{i\in[3]}(\epsilon_{i}+m_{i}-p_{i})^{2}$, and can be further written as $0<\sum_{i\in[3]}(m_{i}-p_{i})(m_{i}-p_{i}+2\epsilon_{i})$. Since $-s<2\epsilon_{i}<s$ and $m_{i}-p_{i}$ can only take the values of $\{...,-2s,-s,0,s,2s,...\}$, the two terms must have the same sign or one of them equals 0. The proof is completed.
Implementation Details. We train CelebAMask-HQ and Cityscapes with a batch size of 256 with AdamW optimizer for 80K and 8K steps, respectively. The initial learning rate is set as $2\times 10^{-5}$ and $5\times 10^{-5}$. Linear learning rate scheduler is adopted. For COCO-Stuff dataset, we use a constant learning rate of $1\times 10^{-5}$ with a batch size of 128 for 320K steps. The COCO-Stuff indicates the version of COCO-Stuff-164k.
Appendix B More Empirical Studies and Visual Results
Sampling from Other Distributions. Fig. 8 shows the synthesized images which are sampled with different perturbations $\epsilon\sim\mathrm{U}(-\beta^{\prime},\beta^{\prime})$. It can be observed that the model can obtain high-quality synthetic results only when the perturbations in the inference stage and the training stage follow the same distribution. Specifically, the synthesized images are dim when $\beta^{\prime}<\beta$, and overexposed when $\beta^{\prime}>\beta$.
Figure 8: Image synthesis results sampled from other distributions. We sample with different perturbation $\epsilon\sim\mathrm{U}(-\beta^{\prime},\beta^{\prime})$. In the training stage, the perturbation follows $\mathrm{U}(-\beta,\beta)$.
More Visual Results on One-step Generation. Fig. 9 shows that our model can generate competitive images with only one inference step.
Figure 9: Synthesized images with one inference step.
Selection of ODE solvers. Fig. 10 provides visualization results on different ODE solvers. A stronger solver has the potential to bring about better results.
Figure 10: The influences of ODE solvers. (a) Euler indicates sampling with euler-25 solver. (b) RK45 indicates the Runge-Kutta method of order 5(4).
More Visual Results on Cityscapes. Fig. 11 shows the results of multi-modal image synthesis. The appearance of the synthesized scene varies according to different random seeds. Moreover, our model can learn the laws of shadows and light angles (the third column), demonstrating its great potential.
Fig. 12 shows visual results of semantic segmentation. Our model accurately segments the objects and assigns the correct semantic labels.
Figure 11: Image synthesis results on Cityscapes. We show the results under three random seeds for each semantic mask. The first row: semantic layouts. The second to the fourth row: synthesized results.
Figure 12: Semantic segmentation results on Cityscapes. The color black in the ground truth indicates the ignored regions.
Appendix C Discussion
Limitations. Some previous work, such as Stable Diffusion, SDXL show that scaling up the training data is essential to exploit the potential of diffusion models fully. Due to the limited computational resources and the number of image-mask pairs in the existing dataset, we leave it to future work. In addition, although this paper does not focus on the network structure and the design of ODE solvers, we believe there is room for exploration to further improve performance.
Broader Impacts. Our SemFlow extends the diffusion model from the generative domain to the segmentation task and narrows the gap with traditional discriminative models. It also functions as a unified framework to bind semantic segmentation and image synthesis. We hope this will motivate people to explore the unification of low-level and high-level vision. In terms of social impact, it will encourage artistic content creation, but at the same time may face the problem of fake content.
Safeguards. The content produced by a generative model is highly correlated with its training data. Ensuring fair and clean training data can effectively prevent models from generating harmful content.